OpenStudy (anonymous):

(x2∗e^−x^2)dx integrate

OpenStudy (anonymous):

if this is \[\int x^2e^{-x^2}dx=\int \frac{x^2}{e^{x^2}}dx\] i am fairly sure you will not find a nice closed form for this itegral

OpenStudy (anonymous):

so how do i proceed?